OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y - x2 = -3x - 10? Answer (2, 0) and (-5, 0) (-2, 0) and (5, 0) (-2, 0) and (-5, 0) (2, 0) and (5, 0)

OpenStudy (asnaseer):

this looks like a duplicate of a question I already helped you on: http://openstudy.com/updates/4fae825de4b059b524f9b8be

OpenStudy (anonymous):

(-2, 0) and (5, 0) 0-x^2=-3x-10 x^2-3x-10=0 (x-5)(x+2) x=5, x=-2

OpenStudy (anonymous):

therefore x intercepts (5,0) and (-2,0)